8) Calculate the wavelength in meters of the wave associated with a 1.00kg mass moving at 1.00km/hr?
Answer #1

v = 1 km/hr
= 1*1000 m / 3600 s
= 0.278 m/s

Given:
mass, m = 1 Kg

use:
lambda = h / (m*v)
lambda = (6.626*10^-34) / (1*0.278)
lambda = 2.383*10^-33 m
Answer: 2.38*10^-33 m